

********************************************************************
UARS WEEKLY STATUS REPORT - 04 September 1998
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
SOLSTICE has requested CDHF to try to determine why the "wall-clock"
times of SOLSTICE production runs is steadily increasing with the
mission day number (not the processing date). The SOLSTICE CPU times
(as well as other resources) are constant over the mission, and the
wall-clock time increase occurs only in the Production environment,
and not under RSS services. The wall-clock time increase is signifi-
cant. For example, our Level 2 job step, which uses about 14 minutes
of CPU time, takes about 40 to 50 minutes for early mission days,
but takes 200-300 minutes for recent mission days. If any other
instrument team has experienced this phenomenon, your input might
be helpful to this investigation.
***********
SUSIM
***********
SUSIM has successfully gathered its normal solar, offset, and occulation
data during this past week without apparent problems.
***********
PEM
***********
PEM was not active
GSFC has proposed an OBC operation plan for the activation of PEM boom
instruments during the daylight portions of every orbit. This OBC operation
plan is accepted by PEM. PEM has proposed a series of test procedures for
activation and is waiting response from GSFC. PEM is also waiting for GSFC
scheduling and implementation of the OBC plan and its operation test procedures.
***********
HALOE
***********
HALOE was powered on beginning of day August 28, 1998. All engineering
parameters are nominal and the science data channels are within the
expected operating range.
The HALOE power on was changed to reduce initial power demand caused by the
operational heaters. Because the survival temperatures were lowered to
reduce the battery supplied survival power to zero (only auxiliary array
"daytime" power is used),the operational heater thermostats are already
activated at power on. Since the heater and gimbal motors are on a
separate relay (pulse bus), the power on procedure was changed so that only
the quiet or electronics bus is turned on. The survival bus continues to
provide heating with the survival heaters during the solar array day. When
the instrument temperature is above the operational heater "off" set point,
the survival bus can be turned off and the heater/motor bus enabled. At
that point occultation measurements can begin. This occurs after
approximately 8 hours for a low beta power on. This procedure is not
necessary for "after graze" power on's.
***********
MLS
***********
MLS is scheduled to begin operations on Sep 8. The MLS team will operate the
instrument in a mode for diagnosing the reoccurance of slips in the antenna
scan mechanism, and to attempt to find a method of operating the instrument
which minimizes the impact of these slips on scientific data.
***********
HRDI
***********
HRDI is currently on and operating normally. The current on period is
scheduled to be August 30-September 7.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no other problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 22 August 1998 (Orbit 37951, GMT Day 234) through 28
August 1998 (Orbit 38054, GMT Day 240).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
WINDII 235/2234 37979 Off
HALOE 240/0037 38041 On
The beta angle ranged between the angles of 16.5 degrees to 13.0 degrees
for this report period. The beta angle is now decreasing toward a minimum
angle of 52.0 degrees on 11 September 1998 (DOY 254).
A Forward-to-Reverse Yaw Around maneuver (number 71) was successfully
executed in Orbit 38004 at 237/13:46:24 UTC on 25 August 1998. The SSPP
Instruments (ACRIM, SOLSTICE, and SUSIM) were enabled. Due to the SA
position at 269 degrees, the effective night length of the spacecraft after
the maneuver was approximately 50 minutes. Post maneuver analysis of data
indicate that during the Yaw Around, the end-of-night load bus voltage was
23.8 volts and the average depths of discharge (DODs) on Batteries 2 and 3
were approximately 23.0 and 22.5 percent respectively.
UARS passed through the annular solar eclipse on two successive orbits on
22 August 1998. The first eclipse entry was 234/01:40:11 with an exit of
234/01:54:54 and a maximum coverage of 28.605 %. The eclipse was near
midday and the On Board Computer Power Monitor Processor (PMON) was
disabled to prevent heater load turn on and re-enabled the following orbit.
The second eclipse entry was 234/03:23:33 with an exit of 234/03:30:12 and
a maximum coverage of 3.551 %. The heater loads were operated normally
during this orbit. The next solar eclipse is annular and will occur on 16
February 1999
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 0.9 to 1.3 degrees C. Additional battery performance data
for each battery on 21 August 1998 and 28 August 1998 are provided in
Appendix D.
The clock error ranged from -2.1 to 6.0 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 07
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 00
% HRDI 08
% HALOE 02
% WINDII 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 05
Two (2) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 98/210 - 2287.496941
Transponder B: 98/211 - 2287.499161
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/239 show the observatory data loss to be
21 hours, 04 minutes, 46.100 seconds (an increase of 0.768 seconds since
last report period). This is a 0.0347 percent data loss which equals a
99.9653 percent data capture for the mission.
The increase of 0.768 seconds occurred over two days:
7 0.224 seconds on Day 233 due to line errors, and
7 0.544 seconds on Day 236 due to hits in the data.
The Applications Processor (AP) Operational Pack for software version 10.12
was delivered to the FOT on 24 August 1998. Ground system software testing
was resumed on 25 August 1998 and continued all week. An OBC dump capture
problem was encountered on 27 August and has been tentatively isolated to
the IP configuration with the UARS simulator. Testing continues.
A scheduled outage of the air conditioning unit in Building 14 on 22 August
1998 caused excessive heat buildup resulting in several of the computer
systems located in the Command Management Facility (CMF) area to be taken
off-line and later returned. The FOT experienced difficulty in trying to
transfer the daily stored command (ATC) and ephemeris loads to the
Applications Processor (AP) from DFLX. After several hours of diagnosis,
it was discovered that one of the Network Adapters between CMF and MSOCC
needed to be reset in order for a session between the AP and DFLX to be
established. Once this Network Adapter was reset, the FOT was able to
successfully retrieve daily loads. However, by the time the loads were
retrieved, the FOT had only 3 and 1/2 minutes of TDRS uplink time to get
the necessary loads to the spacecraft before the ATC processor would halt.
The FOT managed the uplinks and interrupted stored command execution was
averted.
FOT and MSOCC DOC operators were unsuccessful in using the backup tape
transfer methodology for CMS data transfer to the Application Processor
during the Building 14 outage on 22 August 1998. The MSOCC System Engineer
investigated the RCMS Load Tape-to-APS transfer process and was successful.
FOT plans on testing this transfer process next week to verify that it
works, and to verify their knowledge of the procedure itself.
FOT scheduled one extra TDRS pass on GMT Day 243 and two extra passes on
Day 244 to accommodate IP testing. These events will be used to test 512
kbps NBTR playbacks into the Tape Processing Facility (TPF) using the IP
data lines.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
FOT continues to improve a power profile for UARS based on the current
status of the Modular Power System.
The potential Narrowband Tape Recorder (NBTR) noisy playback incidents seem
to be one of a kind since subsequent playbacks have been normal. It can
not be determined whether the problem was on the spacecraft or on the
ground. Anomaly Reports (AR) were issued as historical references to be
reviewed in the event noisy playbacks occurs again.
YEAR 2000 OPERABILITY SUPPORT
UARS Test and Training Simulator (UTTS) No change to report.
User Planning System (UPS) - No change to report.
Multi Satellite Operations Control Center (MSOCC) - No change to report.
Command Management System (CMS) - The FOT attended the weekly CMS Year 2K
Testing status on 27 August. Topic of discussion was the general status of
the testing.
The back-up CMS system (UCMSB) appears to have all of the necessary system
and operational support files to accommodate Year 2K testing. However,
until the FOT receives a complete set of planning aid files from FDF, not
much testing can be conducted. In the meantime, the FOT copied several
1998 FDF planning aid files to UCMSB and processed them as a general test
of the CMS software on the back-up system. These 1998 files seemed to
process fine.
FOT received a partial set of Year 2K planning aid files from Flight
Dynamics Facility (FDF) on 26 August 1998. Due to network configuration
problems with UCMSB, these test files were sent via FORMATS to a special
test directory on the prime CMS system (UCMS). From there, the test files
were copied to UCMSB were they were converted from .IBM files to .RDB using
the FORMATS poller on the back-up CMS system. Actual CMS processing of
these planning aids is expected to take place early next week.
The FOT met with CMS Year 2K test team representative, Sidney Kirschner, to
further discuss the test plan for Year 2K testing on the back-up system.
The FOT identified necessary input and output files for CMS processing and
also provided information regarding the types of TDRS contacts it utilizes.
This latter information is necessary because Sidney is requesting that UPS
development team leader, John Brown, generated a few sets of Year 2K TDRS
contacts to accommodate testing on UCMSB.
Flight Dynamics Facility (FDF) - Year 2K planning aids have been generated.
OBC PMON FLIGHT SOFTWARE UPDATE
Flight Software - The Flight Software is complete.
Ground Support Software - Complete. The Applications Processor operational
disk pack has been delivered for FOT testing.
Verification Testing - Testing of the operations pack compilation of the
software has been initiated.
IP NETWORK TESTING
Space Network (SN) - I/P testing was conducted during TDRSS contacts at
239/13:30 to 239/13:38 and at 239/21:31 to 239/21:37. GCMRs and ODMs were
successful. Q-channel stored Command Load and ephemeris dumps were also
successful.
Testing with the NCC 98 system to support routine flight operations is
being planned.
Deep Space Network (DSN) and Ground Spaceflight Tracking and Data Network
(GSTDN) - Working with the DSN test lead for I/P testing to verify UARS
emergency backup sites.
UARS PROPULSION MODULE B/D THRUSTER LINE FILL
The B/D thruster refill effort continues. Jackson and Tull Jackson have
ordered parts for the new electronics. Earliest projection of Latch Valve
test at GSFC are late September 1998.
UARS RE-ENGINEERING - TELEMETRY PROCESSING
Flight operations is providing support to the re-engineering effort to
process the UARS telemetry on the TOMS/EP control center computer system.
UARS COMMAND RE-HOST PHASE 1 DEMONSTRATION
Flight operations continues to support the update of the Lockheed Martin
SCS-21 control center system to support a phase 1 feasibility demonstration
The SCS-21 equipment has been assembled and testing has been initiated.
Testing of SCS-21 with Network Control Center 98 (NCC 98) system and the
UARS Test and Training Simulator (UTTS) for NCC data and real time
commanding was attempted. Testing continues and minor correction
implemented.
MEETINGS
A Project Configuration Control Board was conducted on 28 August 1998 to
review and approve Version 63 of the On Board Computer (OBC) Flight
Software dated 01 June 1998. OBC Flight Software Version 63 will be
uplinked following successful verification of the corresponding ground
system software operational pack currently under test.
No UARS Battery telecon was held this week. The next telecon has been
scheduled for 25 August 1998 at 1:00 pm.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is scheduled for scheduled for 29
September 1998.
The next annular solar eclipse will occur on 16 February 1999 over the
Indian Ocean, Australia and the South Pacific Ocean.
PERSONNEL CHANGES
None
Attachments: Appendix A Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 28 August 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
37958 234/1238 98-306 HRDI HTR Cycle to 100 amp/min
37964 234/2111 98-307 HRDI HTR Cycle to 125 amp/min
37967 235/0320 98-308 HRDI HTR Cycle to 150 amp/min
37979 235/2234 98-309 WINDII Power OFF
37997 237/0329 98-310 HRDI HTR Cycle to 175 amp/min
38002 237/1118 98-311 HRDI Htr Cycle to 125 amp-min
38009 237/2243 98-312 HRDI Htr Cycle to 150 amp-min
38040 239/2332 98-314 HRDI Htr to 50 Amp mins
38041 240/0037 98-313 HALOE Power ON (with PB off)
38045 240/0820 98-315 HALOE Pulse Bus on
38052 240/1937 98-316 HRDI Htr Cycle to 100 amp min
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 28 August 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRIFESURVTEMP Y-HI HRDI IFE SURVIVAL TEMP 37950
QTSISAIF Y-HI SSPP Baseplate by SISA I/F 37972-38054
HRPSEQBIMON Y-HI HRDI PSE Quiet Bus Cur. Mon. 37972-38014
WIEUTEMP Y-LO WINDII EU Temperature 37985-38054
SOSTMPRNTRN Y-LO Temperature Entrance 38001-9
ACTEMPANA2 Y-LO Temp Sensor Electronics 38002-9
HAH28VA R-LO HALO28V HTRMTR 38041-45
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 28 August 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
-No clock adjusts for this report period
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 28 August 1998
21 Aug 1998 - GMT Day 233 Beta = 16.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-3.99 +0.0/ +0.0 18.1 0.000 0.0
2 +67.2/-0.0 +7.09/+4.83 +31.2/-16.4 23.8 1.023 23.1
3 +39.2/-33.6 +5.79/+3.87 +30.4/-16.8 23.8 1.023 22.4
28 Aug 1998 - GMT Day 240 Beta = 13.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -2.73/-4.31 +0.0/ +0.0 18.1 0.000 0.0
2 +89.6/ 0.0 +7.09/+4.51 +31.2/-15.2 24.2 1.020 23.1
3 +78.4/ 0.0 +5.79/+3.55 +30.4/-16.0 24.2 1.020 22.6
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes (V/T 5)
APPENDIX E
UARS ANOMALY REPORTS
Week ending 28 August 1998
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-052 FOT N/A 234/1600 N/A Unable to Pull Loads
98-045 FOT 37973 235/1210 14170 Generic Late Acquisition
98-045 FOT 38000 237/0735 14170 Generic Late Acquisition
98-053 FOT 38030 239/0813 20167 Late Acq. Reason Unk.
98-054 FOT 38045 240/0820 N/A AP Crashed, ODN down
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-047 FOT 37856 227/1713 20149 NOISY NBR P/B
98-053 FOT 38030 239/0813 20167 LATE ACQ.
98-054 FOT 38045 240/0820 ODN SYSTEM HUNG
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-045 FOT See list Aug 1998 14170 GEWNERIC LATE ACQ
98-052 FOT 234/1559 GRD LOAD TRANSFER